# Nonlocal mass creation by metricized accelerations in Euclidean matterspace

### v1

@@ 4 Shannon information in the yin-yang balance of distributed energies
The stabilizing balance of internal counter-forces in the extended charge was argued by Poincaré [13]. Our physical interpretation of the contracted Bianchi identities (5) through the dialectical unity and struggle of mechanical repulsions and attractions corresponds not only to Poincaré’s guidance, but also agrees with the stable cosmology of Kant. The latter was supported by the Hegelian ideas for the primacy of repulsion in the visible existence of volumetric matter (without the consequences of gravitational collapse or the Beantley paradox). Repulsion provides matter for attraction, as Hegel explained from the dialectical stability of material profiles on earth and in the cosmos. Below we mathematically decipher the Hehelean balance of repulsion and attraction in terms of positive (yang, active) and negative (yin, reactive) energies in the Shannon information continuum.
@@
The Lagrange equation of motion (6) for the temporal axis, $\mu\Rightarrow 0$, in the flat-space physics (4) with $B^{i}_{o}=0$ and $B^{o}_{o}=B^{i}_{i}=B/2$,
@@
$\cases{\nabla_{o}\frac{B}{2}\equiv\partial_{o}\frac{B}{2}=\nabla_{o}B^{o}_{o}+\nabla_{i}B^{i}_{o}\equiv[\partial_{o}B^{o}_{o}+\Gamma_{o\lambda}^{o}B^{\lambda}_{o}-\Gamma_{oo}^{\lambda}B_{\lambda}^{o}]\cr+[\partial_{i}B^{i}_{o}+\Gamma_{i\lambda}^{i}B^{\lambda}_{o}-\Gamma_{io}^{\lambda}B_{\lambda}^{i}]=\partial_{o}\frac{B}{2}+[\Gamma_{io}^{i}B^{o}_{o}-\Gamma_{io}^{j}B^{i}_{j}],\cr\Gamma_{io}^{j}[\nabla^{i}a_{j}+\nabla_{j}a^{i}]=\Gamma_{io}^{i}B^{o}_{o}=\Gamma_{io}^{i}B^{j}_{j}=2\Gamma_{io}^{i}\nabla_{j}a^{j},}$ (10)
can be read as the bounds for admissible 3-potentials $g_{i}\equiv g_{oi}/g_{o}$ in the relativistic condition $\Gamma_{io}^{j}B^{i}_{j}=\Gamma_{io}^{i}B/2$. Recall from (5) that $\Gamma_{io}^{j}=\Gamma_{io}^{i}=0$, if $g_{1}=g_{2}=g_{3}=0.$
@@
Geometrodynamic equations (6)-(7) for three spatial axes, when $\mu\Rightarrow i$, can be simplified and solved analytically for the stationary self-organization of multi-vertex matterspace with $g_{o}(x)=g_{o}({\bm{x}}),g_{j}=0$, $a_{i}(\bm{x})=-\partial_{i}lng_{o}(\bm{x})=\partial_{i}ln[1/g_{o}(\bm{x})],a^{j}(\bm{x})=\delta^{jk}\partial_{k}lng_{o}(\bm{x})\equiv\partial^{j}lng_{o}(\bm{x})$:
@@
$\cases{\nabla_{\nu}T^{\nu}_{i}=\frac{1}{g_{o}}{\partial_{j}(g_{o}T^{j}_{i})}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{8\pi c^{2}g_{o}}\{\partial_{i}[g_{o}(2\partial_{j}a^{j})]-\partial_{j}[g_{o}(\partial_{i}a^{j}-\delta^{jk}\partial_{k}a_{i})]\}=0,\cr\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}[(\partial_{i}lng_{o})\partial_{j}\partial^{j}lng_{o}+\partial_{i}\partial_{j}\partial^{j}lng_{o}]=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}[(\partial_{j}lng_{o})\partial^{j}\partial_{i}lng_{o}+\partial_{j}\partial^{j}\partial_{i}lng_{o}],\cr\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\left(\partial_{i}ln\frac{1}{g_{o}(\bm{x})}\right)\partial_{j}\partial^{j}ln\frac{1}{g_{o}(\bm{x})}=\frac{\varphi_{\!{}_{{}_{G}}}^{2}}{4\pi}\left(\partial_{j}ln\frac{1}{g_{o}(\bm{x})}\right)\partial^{j}\partial_{i}ln\frac{1}{g_{o}(\bm{x})},\cr\varphi_{\!{}_{{}_{G}}}^{2}{a_{i}}\partial_{j}a^{j}=\varphi_{\!{}_{{}_{G}}}^{2}{a_{j}}\partial^{j}a_{i},\ {\sqrt{G}}\mu_{pr}\varphi_{\!{}_{{}_{G}}}a_{i}\equiv f_{i}\equiv-{\sqrt{G}}\mu_{pr}\partial_{i}W,\ W\equiv-{\varphi_{\!{}_{{}_{G}}}}ln(\frac{1}{g_{o}}),\cr\frac{1}{g_{o}(\bm{x})}=1+\frac{r_{1}}{|{\bm{x}}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bm{x}}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bm{x}}-{\bm{\xi}}_{n}|},\ r_{k}\equiv{\sqrt{G}}m_{k}/\varphi_{\!{}_{{}_{G}}},\cr\varphi_{\!{}_{{}_{G}}}\{a_{1};a_{2};a_{3}\}=-\varphi_{\!{}_{{}_{G}}}\frac{\frac{r_{1}({\bm{x}}-{\bm{\xi}}_{1})}{|{\bm{x}}-{\bm{\xi}}_{1}|^{3}}+\frac{r_{2}({\bm{x}}-{\bm{\xi}}_{2})}{|{\bm{x}}-{\bm{\xi}}_{2}|^{3}}+...+\frac{r_{n}({\bm{x}}-{\bm{\xi}}_{n})}{|{\bm{x}}-{\bm{\xi}}_{n}|^{3}}}{1+\frac{r_{1}}{|{\bm{x}}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bm{x}}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bm{x}}-{\bm{\xi}}_{n}|}}\equiv-{\bm{\partial}}W(\bm{x}).}$ (11)
The metrically organized field intensity $\varphi_{\!{}_{{}_{G}}}\!{a_{i}}(\bm{x})$ for the probe charge ${\sqrt{G}}m_{pr}$ depends on the Shannon-type potential $W(\bm{x})=-\varphi_{\!{}_{{}_{G}}}ln[1+\varphi^{-1}_{\!{}_{{}_{G}}}\sum({\sqrt{G}}m_{k}/|{\bm{x}}-{\bm{\xi}}_{k}|)]$ for the optimal information transfer [16]. Shannon information exchange is embedded in the geometric organization of multi-vertex matterspace. Here the positive self-energy of active emergent charges $\varphi_{\!{}_{{}_{G}}}\sum{\sqrt{G}}m_{k}\equiv m_{k}c^{2}>0$ is compensated by the volumetric integral of the negative potential energies ${\sqrt{G}}\mu_{p}({\bf x})W({\bf x})<0$ of the reactive local densities ${\sqrt{G}}\mu_{p}({\bf x})$:
@@
$\cases{{\sqrt{G}}\mu_{p}({\bf x})\equiv\frac{[-{\bm{\partial}}W({\bf x})]^{2}}{4\pi\varphi_{\!{}_{{}_{G}}}}=\frac{{\bm{\partial}}^{2}W({\bf x})}{4\pi}\equiv{{\sqrt{G}}\mu_{a}({\bf x})},\cr\int\!d^{3}x\!{\sqrt{G}}\mu_{a}({\bf x})\!\equiv\!\int\!\frac{\!d^{3}x\!}{4\pi}{\bm{\partial}}^{2}W({\bf x})\!=\!\!\int\!\frac{\!d{\bm{S}}\!}{4\pi}{{\bm{\partial}}W({\bf x})}\!=\!{\varphi}_{\!{}_{{}_{G}}}\!\sum_{k}\!r_{k}\!\equiv\!{\sqrt{G}}\sum_{k}\!m_{k},\cr\int\!\!{\sqrt{G}}\mu_{p}({\bf x})d^{3}x\!\equiv\!\frac{{\varphi}_{\!{}_{{}_{G}}}}{4\pi}\!\!\int\!\!\left(\!\frac{\frac{{({\bf x}-{\bf\xi}_{1})r_{1}}}{|{\bm{x}}-{\bm{\xi}}_{1}|^{3}}+\frac{({\bf x}-{\bm{\xi}}_{2})r_{2}}{|{\bf x}-{\bm{\xi}}_{2}|^{3}}+...+\frac{({\bf x}-{\bm{\xi}}_{n})r_{n}}{|{\bf x}-{\bm{\xi}}_{n}|^{3}}}{1+\frac{r_{1}}{|{\bf x}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bf x}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bf x}-{\bm{\xi}}_{n}|}}\!\right)^{\!2}\!\!\!d^{3}x=\!{\sqrt{G}}\sum_{k}\!m_{k},\cr\int\!\!{\sqrt{G}}\mu_{p}(\!{\bf x}\!)W\!(\!{\bf x}\!)d^{3}x\!\equiv\!-\!\!\int\!\!\frac{d^{3}x}{4\pi}\!\!\left(\!\!\frac{\varphi_{\!{}_{{}_{G}}}\!\!\sum_{k}\!\!\frac{{({\bf x}-{\bm{\xi}}_{k})r_{k}}}{|{\bf x}-{\bm{\xi}}_{k}|^{3}}}{1+\sum_{k}\!\!\frac{r_{k}}{|{\bf x}-{\bm{\xi}}_{k}|}}\!\!\right)^{2}\!\!\!\!ln\!\left(\!\!1\!+\!\sum_{k}\!\!\frac{r_{k}}{|{\bf x}-{\bm{\xi}}_{k}|}\!\!\right)\!=\!-\!c^{2}\!\sum_{k}\!\!m_{k},\cr\int\!d^{3}x[{\sqrt{G}}\mu_{a}({\bf x})\varphi_{\!{}_{{}_{G}}}+{\sqrt{G}}\mu_{p}({\bf x})W\!({\bf x})]=0.}$ (12)
@@
The appearance of active charge densities ${\sqrt{G}}\mu_{a}(x)$ requires metric stresses and local energy costs for the generation of rest energy densities $\mu_{a}(x)c^{2}>0$ due to the positive self-potential $\varphi_{\!{}_{{}_{G}}}\equiv c^{2}/{\sqrt{G}}$. Positive rest energy of macroscopic bodies can be considered as kinetic and related to microscopic motions or fast micro-oscillations. Passive (or reactive) charge density ${\sqrt{G}}\mu_{p}(x)$ is affected by negative potential $W$ for information rates or metric compounds. The local energy imbalance $\mu_{a}(x)c^{2}+{\sqrt{G}}\mu_{p}(x)W(x)\neq 0$, where $\mu_{a}(x)=\mu_{p}(x)$, occurs under the volumetric zeroing in (11) of active (Yang, kinetic) and reactive (Yin, potential) self-energy integrals in the non-local distribution of matterspace with metric stresses and information exchanges.

### v2

@@ 4.1 Static equilibrium of Euclidean matterspace
The stabilizing balance of internal counter-stresses in the extended charge was argued by Poincaré [17]. Our physical interpretation of the contracted Bianchi identities (7) or (8) through the dialectical unity and struggle of kinetic repulsions and metric attractions corresponds not only to Poincaré’s guidance but also agrees with the stable cosmology of Kant. The latter was supported by the Hegelian ideas for the primacy of kinetic repulsion in the visible existence of volumetric matter (without the consequences of gravitational collapse or the Beantley paradox). Repulsion provides matter for Newtonian attraction, as Hegel explained from the dialectical stability of material profiles on Earth and in the cosmos. Below we decipher the Hehelean balance of repulsion and attraction in mathematical terms of positive (yang, active) and negative (yin, reactive) energy densities in the Shannon information continuum with the zero integral (5).
@@
The Lagrange equation of motion (7) for the temporal axis $(\mu\Rightarrow 0)$ in the flat-space physics (4) with $B_{o}^{i}=0,B_{i}^{o}\neq 0$ and $B_{o}^{o}=B_{i}^{i}=B/2>0$ can be expressed as follows:
@@
$\left\{\begin{array}[]{l}\nabla_{o}\frac{B}{2}\equiv\partial_{o}\frac{B}{2}=\nabla_{o}B_{o}^{o}+\nabla_{i}B_{o}^{i}\equiv\left(\partial_{o}B_{o}^{o}+\Gamma_{o\lambda}^{o}B_{o}^{\lambda}-\Gamma_{oo}^{\lambda}B_{\lambda}^{o}\right)\\ +\left[\partial_{i}B_{o}^{i}+\Gamma_{i\lambda}^{i}B_{o}^{\lambda}-\Gamma_{io}^{\lambda}B_{\lambda}^{i}\right]=\left(\partial_{o}\frac{B}{2}-\Gamma_{oo}^{j}B_{j}^{o}\right)+\left[\Gamma_{io}^{i}B_{o}^{o}-\Gamma_{io}^{j}B_{j}^{i}\right],\\ \Gamma_{io}^{i}B_{j}^{j}=\Gamma_{io}^{i}B_{o}^{o}=\Gamma_{io}^{j}B_{j}^{i}+\Gamma_{oo}^{j}B_{j}^{o}\equiv\Gamma_{\nu o}^{j}B_{j}^{\nu}=\Gamma_{\nu o}^{\mu}B_{\mu}^{\nu}\end{array}\right.$ (15)
These dynamical relations can be read as the bounds for admissible 3-potentials $g_{i}\equiv g_{oi}/g_{o}$ in the time-varying condition $\Gamma_{io}^{i}=2\Gamma_{\nu o}^{\mu}B_{\mu}^{\nu}/B$ of the flat matterspace. Recall from (6) that $\Gamma_{io}^{j}=\Gamma_{io}^{i}=0$ in the absence of gyro-potentials $g_{1}=g_{2}=g_{3}=0$, wich leads to $B_{j}^{o}=0$ for such laminar flows or static states without rotations.
@@
Geometrodynamic equations (8) for three spatial axes, when $\mu\Rightarrow i$ in (7), can be simplified for the static equilibrium where one can use $g_{i}=0,g^{oi}=$ $0,\Gamma_{io}^{j}=\Gamma_{ik}^{j}=0,a_{o}\equiv u^{i}f_{io}=0,a^{o}\equiv g^{o\nu}a_{\nu}=0,\nabla_{i}a^{j}=\partial_{i}a^{j},a_{i}=-\partial_{i}\ln g_{o}(\bm{x})=$ $\partial_{i}\ln\left[1/g_{o}(\bm{x})\right],a^{j}=g^{ji}a_{i}=\delta^{ji}\partial_{i}\ln g_{o}(\bm{x})\equiv\partial^{j}\ln g_{o}(\bm{x}):$
@@
$\left\{\begin{array}[]{l}\nabla_{\nu}T_{i}^{\nu}=\frac{1}{g_{o}}\partial_{j}\left(g_{o}T_{i}^{j}\right)\\ =\frac{\varphi_{G}^{2}}{8\pi c^{2}g_{o}}\left\{\partial_{i}\left[g_{o}\left(2\partial_{j}a^{j}\right)\right]-\partial_{j}\left[g_{o}\left(\partial_{i}a^{j}-\partial^{j}a_{i}\right)\right]\right\}=0,\\ \frac{\varphi_{G}^{2}}{4\pi}\left[\left(\partial_{i}\ln g_{o}\right)\partial_{j}\partial^{j}\ln g_{o}+\partial_{i}\partial_{j}\partial^{j}\ln g_{o}\right]\\ =\frac{\varphi_{G}^{2}}{4\pi}\left[\left(\partial_{j}\ln g_{o}\right)\partial^{j}\partial_{i}\ln g_{o}+\partial_{j}\partial^{j}\partial_{i}\ln g_{o}\right],\\ \frac{\varphi_{G}^{2}}{4\pi}\left(\partial_{i}\ln\frac{1}{g_{o}(\bm{x})}\right)\partial_{j}\partial^{j}\ln\frac{1}{g_{o}(\bm{x})}=\frac{\varphi_{G}^{2}}{4\pi}\left(\partial_{j}\ln\frac{1}{g_{o}(\bm{x})}\right)\partial^{j}\partial_{i}\ln\frac{1}{g_{o}(\bm{x})},\\ {\rm or}\ a_{i}\partial^{j}a_{j}=a_{j}\partial^{j}a_{i},\partial^{j}\equiv\delta^{jk}\partial/\partial x^{k}.\end{array}\right.$ (22)
Here we use the mathematical identity $\partial_{i}\partial_{j}\partial^{j}\ln g_{o}(\bm{x})\equiv\partial_{j}\partial^{j}\partial_{i}\ln g_{o}(\bm{x})$ for the partial derivatives of the logarithmic function $\ln g_{o}(\bm{x})$ in Euclidean 3-space $\bm{x}\Rightarrow x^{i}$. The static matterspace with the spherical symmetry has the rectungular coordinate acceleration $c^{2}a_{i}=-Gmx^{i}/\left[r^{2}\left(r+r_{o}\right)\right]$ [16], which corresponds to the non-collapsing equilibrium (11). It turned out that this spatial acceleration quantitatively coincides with the Newtonian field references at large distances from the origin, when $|\mathbf{x}|=r\gg r_{o}\equiv Gm/c^{2}$. But how to find strong field solutions of (11) or self-accelerations $a_{i}$ in corelated matterspace with multiple vertices of mass densities?
4.2 Shannon information references nullify energy integral in equilibrium self-assembly
Now we compare the spherical field potential $W(r)=\varphi_{{}_{G}}\ln g_{o}(r)=-\varphi_{{}_{G}}\ln[1+\varphi_{{}_{G}}^{-1}\left.(\sqrt{G}m/r)\right]$ for the interaction strength $-\partial_{r}W(r)$ exerted on probe (passive) charges $\sqrt{G}m_{pr}$ with the Shannon [19] transmission capacity $C=w\log[1+$ $\left.\left(w_{o}/w\right)\right]\approx-w\log\left(w/w_{o}\right)$, where the noise power is proportional to the information bandwidth $w$. The universal potential $\varphi_{{}_{G}}\equiv c^{2}/\sqrt{G}$ acts as a transmission band for energy potentials $\sqrt{G}m_{k}/r_{k}$ from elementary mechanical charges $\sqrt{G}m_{k}$. One can construct the metric component $g_{oo}\equiv g_{o}^{2}$ and the interaction potential $W(\bm{x})=\varphi_{{}_{G}}\ln g_{o}(\bm{x})$ in any point $\mathbf{x}$ of the continuous matterspace with many information and energy centers in a line of the Shannon theory:
@@ 4.2 Shannon information references nullify energy integral in equilibrium self-assembly
$\begin{array}[]{r}\frac{1}{g_{o}(\bm{x})}=1+\frac{r_{1}}{\left|\bm{x}-\bm{\xi}_{1}\right|}+\frac{r_{2}}{\left|\bm{x}-\bm{\xi}_{2}\right|}+\ldots+\frac{r_{n}}{\left|\bm{x}-\bm{\xi}_{n}\right|},r_{k}\equiv\sqrt{G}m_{k}/\varphi_{{}_{G}},\\ W(\bm{x})=-\varphi_{{}_{G}}\ln\left(1+\frac{1}{\varphi_{{}_{G}}}\sum_{k}\frac{\sqrt{G}m_{k}}{\left|\bm{x}-\bm{\xi}_{k}\right|}\right),\\ \varphi_{{}_{G}}\bm{a}(\bm{x})\equiv-\bm{\partial}W(\bm{x})=-\varphi_{{}_{G}}\frac{\frac{r_{1}\left(\bm{x}-\bm{\xi}_{1}\right)}{\left|\bm{x}-\bm{\xi}_{1}\right|^{3}}+\frac{r_{2}\left(\bm{x}-\bm{\xi}_{2}\right)}{\left|\bm{x}-\bm{\xi}_{2}\right|^{3}}+\ldots+\frac{r_{n}\left(\bm{x}-\bm{\xi}_{n}\right)}{\left|\bm{x}-\bm{\xi}_{n}\right|^{3}}}{1+\frac{r_{1}}{\left|\bm{x}-\bm{\xi}_{1}\right|}+\frac{r_{2}}{\left|\bm{x}-\bm{\xi}_{2}\right|}+\ldots+\frac{r_{n}}{\left|\bm{x}-\bm{\xi}_{n}\right|}}.\end{array}$ (26)
The instantaneous transfer of big data is embedded in a multi-vertex matterspace with the correlated information and mass-energy distribution. The non-locality of macro-organisations is difficult to explain by the empty-space model of dualistic approaches, which deny the continuous matterspace of Cartesian physics.
Now one can relate active and passive mass densities with the Shannon potential $W(\bm{x})$ for energy interactions and information transfer:
$\cases{{\sqrt{G}}\mu_{p}({\bm{x}})\equiv\frac{[-{\bm{\partial}}W({\bf x})]^{2}}{4\pi\varphi_{\!{}_{{}_{G}}}}=\frac{{\bm{\partial}}^{2}W({\bf x})}{4\pi}\equiv{{\sqrt{G}}\mu_{a}({\bf x})},\cr\int\!d^{3}x\!{\sqrt{G}}\mu_{a}({\bf x})\!\equiv\!\int\!\frac{\!d^{3}x\!}{4\pi}{\bm{\partial}}^{2}W({\bf x})\!=\!\!\int\!\frac{\!d{\bm{S}}\!}{4\pi}{{\bm{\partial}}W({\bf x})}\!=\!{\varphi}_{\!{}_{{}_{G}}}\!\sum_{k}\!r_{k}\!\equiv\!{\sqrt{G}}\sum_{k}\!m_{k},\cr\int\!\!{\sqrt{G}}\mu_{p}({\bf x})d^{3}x\!\equiv\!\frac{{\varphi}_{\!{}_{{}_{G}}}}{4\pi}\!\!\int\!\!\left(\!\frac{\frac{{({\bf x}-{\bm{\xi}}_{1})r_{1}}}{|{\bm{x}}-{\bm{\xi}}_{1}|^{3}}+\frac{({\bf x}-{\bm{\xi}}_{2})r_{2}}{|{\bf x}-{\bm{\xi}}_{2}|^{3}}+...+\frac{({\bf x}-{\bm{\xi}}_{n})r_{n}}{|{\bf x}-{\bm{\xi}}_{n}|^{3}}}{1+\frac{r_{1}}{|{\bm{x}}-{\bm{\xi}}_{1}|}+\frac{r_{2}}{|{\bf x}-{\bm{\xi}}_{2}|}+...+\frac{r_{n}}{|{\bm{x}}-{\bm{\xi}}_{n}|}}\!\right)^{\!2}\!\!\!d^{3}x=\!{\sqrt{G}}\sum_{k}\!m_{k},\cr\int\!\!{\sqrt{G}}\mu_{p}(\!{\bf x}\!)W\!(\!{\bm{x}}\!)d^{3}x\!\equiv\!-\!\!\int\!\!\frac{d^{3}x}{4\pi}\!\!\left(\!\!\frac{\varphi_{\!{}_{{}_{G}}}\!\!\sum_{k}\!\!\frac{{({\bf x}-{\bm{\xi}}_{k})r_{k}}}{|{\bm{x}}-{\bm{\xi}}_{k}|^{3}}}{1+\sum_{k}\!\!\frac{r_{k}}{|{\bf x}-{\bm{\xi}}_{k}|}}\!\!\right)^{2}\!\!\!\!ln\!\left(\!\!1\!+\!\sum_{k}\!\!\frac{r_{k}}{|{\bf x}-{\bm{\xi}}_{k}|}\!\!\right)\!=\!-\!c^{2}\!\sum_{k}\!\!m_{k},\cr\int\!d^{3}x[{\sqrt{G}}\mu_{a}({\bm{x}})\varphi_{\!{}_{{}_{G}}}+{\sqrt{G}}\mu_{p}({\bm{x}})W\!({\bm{x}})]=0.}$ (27)
Here the positive rest-energy of active emergent charges $\varphi_{G}\!\sum_{k}\!\sqrt{G}m_{k}\equiv\sum_{k}m_{k}c^{2}$ > 0 is compensated by the volumetric integral of the negative potential energies $\sqrt{G}\mu_{p}(\bm{x})W(\bm{x})<0$ of the passive densities $\sqrt{G}\mu_{p}(\mathbf{x})$ in the Shannon interaction potential from (12). This yin-yang dialectic with the zero energy balance of non-local self-assemblies underlines how informatics can be an integral part of understanding the correlated interplay of distributed energies, forces and big data transmissions within an isolated material system.
@@
The appearance of active charge densities $\sqrt{G}\mu_{a}(x)$ requires metric stresses and local energy costs for the generation of rest energy densities $\mu_{a}(x)c^{2}>0$ due to the positive self-potential $\varphi_{G}\equiv c^{2}/\sqrt{G}$. Positive rest energy of macroscopic bodies can be considered as kinetic and related to microscopic motions or fast micro-oscillations. The passive (or reactive) charge density $\sqrt{G}\mu_{p}(x)$ is affected by the Shannon potential $W(x)<0$ for information rates and metric correlations. The local energy imbalance $\mu_{a}(x)c^{2}+\sqrt{G}\mu_{p}(x)W(x)\neq 0$, where $\mu_{a}(x)=\mu_{p}(x)$, occurs under the volumetric zeroing in (13) of active (yang, kinetic) and reactive (yin, potential) self-energy integrals in the non-local distribution of matterspace with metric stresses and information exchange.
The Yin-Yang worldview has a long history of studies in the East. In Europe, the most consistent proponent of the complete cancellation of the rest energy of a self-gravitating star by its internal binding energies was Pascual Jordan. From the observed coexistence of old and new stars he deduced their spontaneous formation with zero summary energy [20]. At that time, the gravitational binding energy in the strong-field regime was not modelled in general relativity. And the advanced hypothesis of ‘a new and constant production of mass in space’ with zero relativistic energy was quite puzzling.